Find two consecutive odd integers such that the square of the first decreased by the second is 28.

I just don't understand how to set it up and prove that it is right, my guess is obvious 5, -3 Help

Answer by ewatrrr(24785)


Hi
Let x and (x+2)represent the two consecutive odd integers
Question states***
x^2 - (x+2) = 28
x^2 -x -30 = 0
factoring
(x-6)(x+5) = 0
(x-6)= 0 x = 6 |Toss out this solution as it is not an odd number
(x+5) =0 x = -5 the two consecutive odd integers are -5,-3

CHECKING our Answer***
25 -(-3) = 28